# FIRE [Uncaptioned image]: Flexible Integration of Data Quality Ratings for Effective Pretraining

### v2

@@ C.2 Prompt for GPT-4o to compare data quality
C.3 Polynomial spline interpolation for win-rate-percentile function
@@
C.4 Ratings distribution illustration
@@
C.6 Orthogonality-Percentile curve
@@ D.2 Integration Baselines
Following QuRating Wettig et al. (2024), we collect pairwise comparison data and train a reward model based on Sheared-Llama 1.3B Xia et al. (2023). We merge all the evaluation criteria into a single prompt to guide GPT-4o comparison. Below is an illustration of the prompt.
Prompts for Comprehensive Rater Compare two text excerpts and choose the text which 1. has a more polished and beautiful writing style. 2. contains more facts and trivia. Prefer specific facts and obscure trivia over more common knowledge. 3. requires greater expertise and prerequisite knowledge to understand it. 4. has more educational value, e.g., it includes clear explanations, step-by-step reasoning, or questions and answers.
Prompts for Comprehensive Rater Aspects that should NOT influence your judgement: 1. Which language the text is written in 2. The length of the text 3. The order in which the texts are presented Note that the texts are cut off, so you have to infer their contexts. The texts might have similar quality, but you should still make a relative judgement and choose the label of the preferred text. [Option A] {text1} [Option B] {text2} Now you have to choose between either A or B. Respond only with a single word.
@@
D.3 Prompt for Multi-dimension Analysis
@@
Prompts for GPT-4o evaluation You are a data annotation expert. You should judge that {condition} Aspects that should NOT influence your judgement: 1. Which language the text is written in 2. The length of the text 3. The order in which the texts are presented Note that the texts are cut off, so you have to infer their contexts. Here is the text: [TEXT BEGIN] {text} [TEXT END] Please follow the question order to respond. For answer, only respond yes or no. Return the results for each question in the following json format:
Prompts for GPT-4o evaluation [{ "quesion": "Is this text has a polished and beautiful writing style ?", "reason": "Fill in the reason for the judgment here", "answer": "yes/no" }, …]

### v3

@@ C.2 Prompt for GPT-4o to compare data quality
C.3 Reliability Analysis of GPT-4o
QuRatingWettig et al. (2024) points out that GPT is more effective at comparing the relative quality between two data samples than performing absolute quality evaluation. To further assess the reliability of GPT-4o, we use QuRating (Educational Value) as a case study and conduct a win-rate evaluation involving human experts. Specifically, we compare the win rates assigned by human annotators and GPT-4o across different rating percentiles, with results presented in descending order of percentile rating in Table 4. The two sets of win rates exhibit a Pearson correlation of 0.99, indicating strong agreement and suggesting that GPT-4o does not introduce significant bias in the annotation process. This high consistency supports the reliability of using GPT-4o for quality assessment. Furthermore, GPT-4o is used only to estimate win rates between rater-selected samples and random subsets, rather than to assign absolute scores, which further reduces bias.
Percentile 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%
GPT-4o 0.773 0.705 0.625 0.600 0.545 0.513 0.480 0.425 0.340 0.273
Human Expert 0.797 0.698 0.657 0.616 0.533 0.501 0.469 0.433 0.317 0.300
Table 4: Win-rate comparison between GPT-4o and human experts for QuRating (Educational Value) across different rating percentiles.
C.4 Polynomial spline interpolation for win-rate-percentile function
@@
C.5 Ratings distribution illustration
@@
C.7 Orthogonality-Percentile curve
@@ D.2 Integration Baselines
Prompts for Comprehensive Rater Compare two text excerpts and choose the text which 1. has a more polished and beautiful writing style. 2. contains more facts and trivia. Prefer specific facts and obscure trivia over more common knowledge. 3. requires greater expertise and prerequisite knowledge to understand it. 4. has more educational value, e.g., it includes clear explanations, step-by-step reasoning, or questions and answers. Aspects that should NOT influence your judgement: 1. Which language the text is written in 2. The length of the text 3. The order in which the texts are presented Note that the texts are cut off, so you have to infer their contexts. The texts might have similar quality, but you should still make a relative judgement and choose the label of the preferred text. [Option A] {text1} [Option B] {text2} Now you have to choose between either A or B. Respond only with a single word.
@@
Following QuRating Wettig et al. (2024), we collect pairwise comparison data and train a reward model based on Sheared-Llama 1.3B Xia et al. (2023). We merge all the evaluation criteria into a single prompt to guide GPT-4o comparison. Above is an illustration of the prompt.
@@
D.3 Analysis of Computational Cost
This section analyzes the computational cost (measured in FLOPs) for various data rating and integration methods. The focus is on the cost introduced by applying raters to the pre-training data, excluding the main model’s training cost. FIRE achieves more effective integration without introducing significant extra cost beyond the base raters. The FLOPs comparison is summarized in Table 6.
The analysis includes:
  1. 1.
    Rater training (if applicable),
  2. 2.
    Rater inference over the entire dataset,
  3. 3.
    Additional computation for FIRE’s win-rate-based integration.
Method FLOPs
QuRating (single rater) $\approx 8.17\times 10^{20}$
QuRating (mix of criteria, 4 raters) $\approx 3.26\times 10^{21}$
FIRE (mix of criteria, 4 raters) $\approx 3.26\times 10^{21}$
Table 6: Comparison of computational cost (FLOPs) across different rating methods.
QuRating (Single Rater).
Each QuRating model is a 1.3B-parameter transformer. It is first fine-tuned on 500K examples (512 tokens each). The total FLOPs for training this rater is approximately:
$\text{FLOPs}_{\text{train}}\approx 6\cdot N\cdot d^{2}\cdot L\cdot T\approx 4.2\times 10^{19}$
Rater inference over the entire 627B-token dataset consumes:
$\text{FLOPs}_{\text{infer}}\approx 6\cdot 627\times 10^{9}\cdot d^{2}\cdot L\approx 8.13\times 10^{20}$
$\text{FLOPs}_{\text{total}}\approx 8.17\times 10^{20}$
QuRating (Mix of Criteria).
Integrating four QuRating raters (each covering a distinct aspect) requires four forward passes across the full dataset. Since retraining is not needed, total cost is:
$\text{FLOPs}_{\text{mix}}\approx 4\times\text{FLOPs}_{\text{infer}}\approx 3.26\times 10^{21}$
FIRE.
FIRE uses the same four QuRating raters as input, and hence shares the same inference cost. In addition, it estimates the win-rate-percentile mapping using only 20K pairwise comparisons per rater. The additional cost of this step is negligible compared to inference over 627B tokens.
D.4 Integration of Other Raters
FIRE has extremely high scalability: our approach is not directly related to the specific attributes of the raters. Therefore, as long as the raters can provide ratings for the samples and are not completely related to each other, they can be integrated using the FIRE method. We also attempt to integrate another raters: QuRating(Required Expertise), QuRating(Facts and Trivia), DSIR-Book, and DSIR-Wiki. Tabel 7 shows that the average performance of FIRE, after integrating these four raters, is better than each individual rater and the Random baseline, indicating that our integration method is still effective on other raters.
Method ARC-E ARC-C SciQ LogiQA BoolQ HellaSw. PIQA W.G. AVG.
Random 48.2 22.3 84.5 19.7 60.8 32.1 63.5 49.2 47.5
DSIR with Book 36.2 19.5 73.4 21.4 61.8 29.5 62.5 53.6 44.7
DSIR with Wiki 37.2 18.0 76.4 23.0 58.0 27.9 57.3 51.1 43.6
QuRating (R.E.) 50.6 23.2 83.9 22.6 61.4 30.2 59.8 49.8 47.7
QuRating (F.T.) 54.1 23.0 83.5 22.0 60.9 30.4 59.5 51.7 48.1
FIRE 57.1 25.9 84.9 20.8 61.5 31.5 60.1 50.3 49.0
Table 7: Performance comparison by applying FIRE to integrate four raters: QuRating (Required Expertise), QuRating (Facts and Trivia), DSIR-Book, and DSIR-Wiki.
D.5 Prompt for Multi-dimension Analysis
Prompts for GPT-4o evaluation You are a data annotation expert. You should judge that {condition} Aspects that should NOT influence your judgement: 1. Which language the text is written in 2. The length of the text 3. The order in which the texts are presented Note that the texts are cut off, so you have to infer their contexts. Here is the text: [TEXT BEGIN] {text} [TEXT END] Please follow the question order to respond. For answer, only respond yes or no. Return the results for each question in the following json format: [{ "quesion": "Is this text has a polished and beautiful writing style ?", "reason": "Fill in the reason for the judgment here", "answer": "yes/no" }, …]